The difference of two numbers is 49. The second is 8 less then 4 times the first. What are the two numbers.
Answer by elima(1433)

The difference of two numbers is 49. The second is 8 less then 4 times the first. What are the two numbers.
x-y=49
y=4x-8
----------
x-4x-8=49
-3x=57
x=-19
----------
-19-y=49
-y=49+19
-y=68
y=-68
----------
:)